# Online Optimization for Learning to Communicate over Time-Correlated Channels

### v1

@@ III-C Online Optimization Procedure
As mentioned above, this paper models the problem of learning the channel decoder $\mathbf{G}_{t}$ or codebook $C_{t}=\{\mathbf{x}_{t}^{j}\}_{j\in[M]}$ as an online optimization problem. Assume that there is a transmitter-receiver pair, and the transmitter sends codewords to the receiver over the time-correlated channel. Let $\mathcal{D}$ be a feasible set and $\ell:\mathcal{D}\rightarrow\mathbb{R}$ be a loss function. In general, at round $t\in[T]$, the transceiver carries out the following step:
@@
    The receiver makes a decision $D_{t}\in\mathcal{D}$, which can be either a channel decoder $\mathbf{G}_{t}$ or a codebook $C_{t}$. Following [9], we also assume that the receiver can transmit the learned codebook $C_{t}$ to the transmitter via the reliable feedback link.
@@
    The transmitter then sends codewords to the receiver over the time-correlated channel, and then the receiver calculates the symbol error rate based on the received channel output as the loss function $\ell_{t}(D_{t})$.
@@ IV-A Preliminaries
Assumption 1 (Bounded Channel Gain and Channel Noise)
The channel gain is bounded by $\gamma_{H}$ and the additive Gaussian noise is bounded by $\gamma_{W}$ at each round, i.e., for any $t\in[T]$, we have $\mathbb{E}\|\mathbf{H}_{t}\|_{F}\leq\gamma_{H}$ and $\mathbb{E}\|W_{t}\|_{2}=\sigma_{W}\leq\gamma_{W}$.
Based on this assumption, we find that channel outputs $\{\mathbf{y}_{t}^{j}\}_{j\in[M],t\in[T]}$ are bounded by $\sqrt{2(\gamma_{X}\gamma_{H})^{2}+2(\gamma_{W})^{2}}$ and we denote it by $L$ in the following.
@@
Based on the above observation, we utilize the following hinge-type loss as the convex surrogate loss function with respect to $\mathbf{G}_{t}$ with the parameter $r_{t}\geq 1$:
@@
$\mbox{Reg}_{T}:=\sum_{t=1}^{T}\Big{(}\tilde{\ell}_{t}(\mathbf{G}_{t})-\tilde{\ell}_{t}(\mathbf{G}^{*})\Big{)},$ (17)
@@ IV-B Optimistic OMD with Euclidean Regularizer
In this section, we set the regularizer as the Euclidean norm [31], i.e., $\psi_{t}(\mathbf{G})=\frac{1}{2\eta_{t}}\|\mathbf{G}\|^{2}_{F}$ with the learning rate
$\eta_{t}=\frac{D}{\sqrt{1+\sum_{\tau=1}^{t-1}\|\nabla\tilde{\ell}_{\tau}(\mathbf{G}_{\tau})-\mathbf{M}_{\tau}\|_{F}^{2}}}.$ (19)
@@ 5       The receiver calculates ~$\tilde{\ell}_{t}(\mathbf{G}_{t})$;
Under Assumptions 1, if we select the parameters $r_{t}=2d^{*}DL+\frac{1}{\sqrt{T}}$ with the maximum codewords distance $d^{*}:=\max_{i\neq j}\|\mathbf{x}^{i}-\mathbf{x}^{j}\|_{2}\geq\frac{1}{2DL}$, we have
$\frac{1}{T}\sum_{t\in[T]}\mathbb{P}(\mathbf{G}_{t})-\frac{1}{T}\sum_{t\in[T]}\mathbb{E}[\tilde{\ell}_{t}(\mathbf{G}^{*})]$ (21)
$=\mathcal{O}\Big{(}\frac{1}{T}\sqrt{M\sum_{t\in[T]}\sum_{j\in[M]}\mathbb{E}\|\mathbf{y}_{t}^{j}-\mathbf{y}_{t-1}^{j}\|_{2}^{2}}\Big{)}.$
@@
Theorem 1 provides a performance guarantee of the learned channel decoder $\{\mathbf{G}_{t}\}_{t\in[T]}$: if the considered online learning problem is realizable, i.e., $\exists\mathbf{G}^{*}\in\mathcal{G}$ such that the term $\sum_{t\in[T]}\mathbb{E}[\tilde{\ell}_{t}(\mathbf{G}^{*})]$ equals to zero [20, 28], the average expected error probability satisfies $\frac{1}{T}\sum_{t\in[T]}\mathbb{P}(\mathbf{G}_{t})\rightarrow 0$ as $T\rightarrow\infty$. Besides, the term $\|\mathbf{y}_{t}^{j}-\mathbf{y}_{t-1}^{j}\|_{2}^{2}$ reflects the variation of channel outputs across successive rounds transmitted by the same codeword. We observe that as this term decreases, the derived upper bound becomes tighter, indicating that our approach leverages the distribution dependence of $\{\mathbf{H}_{t}\}_{t\in[T]}$ to enhance the efficacy of $\mathbf{G}_{t}$.
@@
$\frac{1}{T}\sum_{t\in[T]}\mathbb{P}(\mathbf{G}_{t})-\frac{1}{T}\sum_{t\in[T]}\mathbb{E}[\tilde{\ell}_{t}(\mathbf{G}^{*})]$ (22)
$=\mathcal{O}\Big{(}\frac{M}{T}\sqrt{(\sigma^{\mathbf{H}}_{1:T})^{2}}+\frac{M}{T}\sqrt{(\Sigma^{\mathbf{H}}_{1:T})^{2}}+\frac{M\sqrt{T}}{T}\sqrt{\sigma_{W}^{2}}\Big{)},$
@@
Corollary 1 implies that the variance $\sigma^{2}_{W}$ of the channel noise $W$, the cumulative variance $(\sigma^{\mathbf{H}}_{1:T})^{2}$ and the cumulative variation $(\Sigma^{\mathbf{H}}_{1:T})^{2}$ of channel gain $\{\mathbf{H}_{t}\}_{t\in[T]}$ can deteriorate the performance of the learned channel decoder $\{\mathbf{G}_{t}\}_{t\in[T]}$. We observe that smaller $(\Sigma^{\mathbf{H}}_{1:T})^{2}$ leads to a tighter upper bound of average expected error probability $\frac{1}{T}\sum_{t\in[T]}\mathbb{P}(\mathbf{G}_{t})$, reflecting that the proposed method utilizes the distribution dependence of the time-correlated channel gain $\{\mathbf{H}_{t}\}_{t\in[T]}$ to enhance the online optimization process.
IV-C Extension: Learning to Communicate over Independently and Identically Distributed Fading Channel
In this subsection, we focus on the I.I.D. fading channel, that is, the channel gains $\mathbf{H}_{1},\mathbf{H}_{2},...,\mathbf{H}_{t}$ are independently sampled from the distribution $F_{\mathbf{H}}$. Consequently, the variation of $\{\mathbf{H}_{t}\}_{t\in[T]}$ between different rounds satisfies $(\Sigma^{\mathbf{H}}_{1:T})^{2}=0$ and the variance of $\mathbf{H}_{t}$ satisfies $(\sigma_{\mathbf{H}_{t}})^{2}=\sigma_{\mathbf{H}}^{2},\forall t\in[T]$.
Hence, the online optimization problem considered before is now converted into a stochastic optimization problem:
$\min_{\mathbf{G}\in\mathcal{G}}$~$\Big{\{}\tilde{\ell}(\mathbf{G})$ (23)
$:=\frac{1}{M}\sum_{j=1}^{M}\sum^{M}_{j^{\prime}\neq j}\big{[}r-\|\mathbf{x}^{j^{\prime}}-\mathbf{G}\mathbf{y}^{j}\|^{2}_{2}+\|\mathbf{x}^{j}-\mathbf{G}\mathbf{y}^{j}\|^{2}_{2}\big{]}_{+}\Big{\}},$
where $r>1$ and $\mathbf{y}^{j}:=\mathbf{H}\mathbf{x}^{j}+W,\forall j\in[M]$. $\mathbf{H}\sim F_{\mathbf{H}}$ is an I.I.D. channel gain with unknown distribution, and $W$ is an I.I.D. additive white Gaussian noise.
Next, we employ the well-established online-to-batch conversion method [20, 21] to solve the problem defined in Eq. (23). Furthermore, we can undertake the generalization analysis of the channel decoder $\bar{\mathbf{G}}$ learned by this way for I.I.D. fading channels. The fundamental idea of this approach is to utilize the I.I.D. channel data in a sequential manner, and execute the online optimization algorithm to minimize the regret. Upon completion of this sequential learning process, the average decision serves as the ultimate output of this online-to-batch conversion procedure.
Input: An online optimization algorithm $\mathcal{A}$
1 for round $t\in[T]$  do
2       Let $\mathbf{G}_{t}$ be the output of algorithm $\mathcal{A}$ for this round;
3       Feed algorithm $\mathcal{A}$ with the loss function ~$\tilde{\ell}_{t}(\mathbf{G}_{t})$ calculated by data $(\mathbf{H}_{t},W_{t})$;
4
5 end for
Return: $\bar{\mathbf{G}}=\frac{1}{T}\sum_{t\in[T]}\mathbf{G}_{t}$
Algorithm 2 Online-to-Batch Conversion
The online-to-batch conversion is performed as illustrated in Algorithm 2. If the online optimization algorithm $\mathcal{A}$ is selected as the Algorithm 1, we can derive the expected error probability of $\bar{\mathbf{G}}$ constructed by $\mathcal{A}$ below.
Theorem 2
If the loss function ~$\tilde{\ell}(\bar{\mathbf{G}})$ is convex w.r.t. $\bar{\mathbf{G}}$, the channel gain $\{\mathbf{H}_{t}\}_{t\in[T]}$ are i.i.d. sampled from $F_{\mathbf{H}}$ and the channel noise $\{W_{t}\}_{t\in[T]}$ are i.i.d. sampled from $F_{W}$, the expected error probability of $\bar{\mathbf{G}}$ satisfies
$\mathbb{P}(\bar{\mathbf{G}})-\min_{\mathbf{G}\in\mathcal{G}}\mathbb{E}_{\mathbf{H},W}[\tilde{\ell}(\mathbf{G})]=\mathcal{O}\Big{(}M\sqrt{\frac{\sigma_{\mathbf{H}}^{2}}{T}}+M\sqrt{\frac{\sigma_{W}^{2}}{T}}\Big{)},$ (24)
where $\mathbb{E}_{\mathbf{H},W}[\tilde{\ell}(\mathbf{G})]$ is the expected risk.
Remark 3
Theorem 2 implies that the expected error probability of $\bar{\mathbf{G}}$ only depends on the variance $\sigma_{\mathbf{H}}^{2}$ of channel gain $\mathbf{H}$ and the variance $\sigma_{W}^{2}$ of channel noise $W$, since $(\Sigma^{\mathbf{H}}_{1:T})^{2}$ reflecting the variation of $\{\mathbf{H}_{t}\}_{t\in[T]}$ becomes zero when channel gain is I.I.D. with the distribution $F_{\mathbf{H}}$. Similarly, if the statistical learning problem defined in Eq. (23) is realizable, i.e., $\exists\mathbf{G}^{*}\in\mathcal{G}$ such that $\mathbb{E}_{\mathbf{H},W}[\tilde{\ell}(\mathbf{G}^{*})]$ equals to zero [12, 13], we have $\mathbb{P}(\bar{\mathbf{G}})\rightarrow 0$, as $T\rightarrow\infty$, providing the theoretical performance guarantee for the proposed online-to-batch conversion method. Hence, $T$ plays a role akin to sample complexity within statistical learning theory [13].
@@ V-A Preliminaries
One potential approach to tackle this challenge involves selecting codebooks for practical transmission from a pre-defined super-codebook $\mathbf{C}=\{C_{i}\}_{i=1}^{N}$ comprising $|\mathbf{C}|=N$ codebooks [9, 37]. This super-codebook can be statically constructed (such as a grid or a lattice [38]) in advance. To adapt to the changing statistical property of the channel noise, we dynamically choose one of these codebooks used for transmitting codewords in each round. Specifically, we select the codebook based on the symbol error rate of codebooks over the channel noise $Z_{t},\forall t\in[T]$. This approach draws inspiration from the Gibbs-algorithm-based codebook expurgation proposed in [9], and is amenable to practical deployment following the paradigm of Adaptive Modulation and Coding (AMC) widely deployed in communication systems [32, 33].
@@ 1 Initialize: $\mathbf{w}^{\prime}_{1}=\arg\min_{\mathbf{w}\in\Omega}\psi_{1}(\mathbf{w})$ ;
$\frac{1}{T}\sum_{t\in[T]}\mathbb{P}(C_{t})-\frac{1}{T}\sum_{t\in[T]}\mathbb{E}[\ell_{t}(C^{*})]$ (28)
$=\tilde{\mathcal{O}}\Big{(}\frac{1}{T}\sqrt{\sum_{i\in[N]}\sum_{t\in[T]}\frac{1}{M}\sum_{j\in[M]}\big{|}\mathbb{P}^{j}_{t}(i)-\mathbb{P}^{j}_{t-1}(i)\big{|}}$
$\quad+\frac{1}{T}\sqrt{\sum_{i\in[N]}\sum_{t\in[T]}\frac{1}{M}\sum_{j\in[M]}\sigma[\mathbbm{1}_{t}^{j}(i)]}\Big{)},$
where $\mathbbm{1}_{t}^{j}(i):=\mathbbm{1}\{\|\mathbf{x}_{t}^{j_{t}}-\mathbf{y}_{t}^{j}\|_{2}\leq\|\mathbf{x}_{t}^{j}-\mathbf{y}_{t}^{j}\|_{2}|\mathbf{a}_{t}=\mathbf{e}_{i}\}$ denotes the indicator of misclassifying the $j$-th codeword when the $i$-th codebook is selected in round $t$. $\mathbb{P}^{j}_{t}(i):=\mathbb{E}[\mathbbm{1}_{t}^{j}(i)]$ denotes the expectation of $\mathbbm{1}_{t}^{j}(i)$, and $\sigma[\mathbbm{1}_{t}^{j}(i)]:=\sqrt{\mathbb{E}[\mathbbm{1}_{t}^{j}(i)-\mathbb{P}^{j}_{t}(i)]^{2}}$ denotes the standard deviation of $\mathbbm{1}_{t}^{j}(i)$.
@@
Theorem 3 presents the theoretical performance guarantee of the learned codebook $\{C_{t}\}_{t\in[T]}$: if $\exists C^{*}\in\mathbf{C}$ such that $\sum_{t\in[T]}\ell_{t}(C^{*})$ is sub-linear, i.e., $\sum_{t\in[T]}\ell_{t}(C^{*})=\mathcal{O}(T^{\alpha}),\alpha<1$, then we have $\frac{1}{T}\sum_{t\in[T]}\mathbb{P}(C_{t})\rightarrow 0$, as $T\rightarrow\infty$. In addition, this theorem demonstrates that the performance of $C_{t}$ becomes better if the expected error probability $\mathbb{P}_{t}^{j}(i)$ of the $j$-th codebword in the same codebook $C_{i},\forall i\in[N]$ varies slowly between successive rounds, i.e., the difference $\big{|}\mathbb{P}^{j}_{t}(i)-\mathbb{P}^{j}_{t-1}(i)\big{|}$ is small. This finding reflects that we can learn the optimal codebook $C_{t}$ if the statistical property of $Z_{t}$ does not change heavily, which indicates that the proposed method leverages the distribution dependence of $\{Z_{t}\}_{t\in[T]}$ to improve the bandit learning process.

### v3

@@ III-C Online Optimization Procedure
As mentioned above, this paper models the problem of learning the channel decoder $\mathbf{G}_{t}$ or codebook $C_{t}=\{\mathbf{x}_{t}^{j}\}_{j\in[M]}$ (a constellation) as an online optimization problem. Assume that there is a transmitter-receiver pair, and the transmitter sends a package containing both the deterministic training sequence including $M$ codewords and random data symbols to the receiver over the time-correlated channel. Let $\mathcal{D}$ be a feasible set and $\ell:\mathcal{D}\rightarrow\mathbb{R}$ be a loss function. In general, at round $t\in[T]$, the transceiver carries out the following step:
@@
    The receiver makes a decision $D_{t}\in\mathcal{D}$, which can be either a channel decoder $\mathbf{G}_{t}$ or a codebook $C_{t}$. Following [8], we also assume that the receiver can transmit the learned codebook $C_{t}$ to the transmitter via the reliable feedback link given that communication resources of the receiver are sufficient.
@@
    The transmitter then sends codewords to the receiver over the time-correlated channel, and then the receiver calculates the symbol error rate based on received channel outputs transmitted by $M$ codewords in the training sequence as the loss function $\ell_{t}(D_{t})$.
@@ IV-A Preliminaries
Assumption 1 (Channel gain and channel noise with bounded expected norm)
The expected norm of channel gain is bounded by $\gamma_{H}$ and the expected norm of additive Gaussian noise is bounded by $\gamma_{W}$ at each round, i.e., for any $t\in[T]$, we have $\mathbb{E}\|\mathbf{H}_{t}\|_{F}\leq\gamma_{H}$ and $\mathbb{E}\|W_{t}\|_{2}=\sigma_{W}\leq\gamma_{W}$.
Based on this assumption, we find that $\{\mathbf{y}_{t}^{j}\}_{j\in[M],t\in[T]}$ are bounded by $\sqrt{2(\gamma_{X}\gamma_{H})^{2}+2(\gamma_{W})^{2}}$ and we denote it by $L$ in the following.
@@
We observe that $r-\|\mathbf{x}^{j^{\prime}}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|^{2}_{2}+\|\mathbf{x}^{j}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|^{2}_{2}=r+2\langle\mathbf{x}^{j^{\prime}}-\mathbf{x}^{j},\mathbf{G}_{t}\mathbf{y}_{t}^{j}\rangle$ is an affine function w.r.t. $\mathbf{G}_{t}$, thus $[r-\|\mathbf{x}^{j^{\prime}}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|^{2}_{2}+\|\mathbf{x}^{j}-\mathbf{G}_{t}\mathbf{y}_{t}^{j}\|^{2}_{2}]_{+}$ is a convex function w.r.t. $\mathbf{G}_{t}$. Based on the above observation, we utilize the following hinge-type loss as the convex surrogate loss function with respect to $\mathbf{G}_{t}$ with the parameter $r_{t}\geq 1$:
@@
$\mbox{Reg}^{\prime}_{T}:=\sum_{t=1}^{T}\Big{(}\tilde{\ell}_{t}(\mathbf{G}_{t})-\tilde{\ell}_{t}(\mathbf{G}^{*})\Big{)},$ (17)
@@ IV-B Optimistic OMD with Euclidean Regularizer
In this section, we set the regularizer as the Euclidean norm [36], i.e., $\psi_{t}(\mathbf{G})=\frac{1}{2\eta_{t}}\|\mathbf{G}\|^{2}_{F}$ with the learning rate $\eta_{t}=\frac{D}{\sqrt{1+\sum_{\tau=1}^{t-1}\|\nabla\tilde{\ell}_{\tau}(\mathbf{G}_{\tau})-\mathbf{M}_{\tau}\|_{F}^{2}}}$.
@@ 5       The receiver calculates ~$\tilde{\ell}_{t}(\mathbf{G}_{t})$;
Under Assumption 1, if we select the parameters $r_{t}=2d^{*}DL+\frac{1}{\sqrt{T}}$ with the maximum codewords distance $d^{*}:=\max_{i\neq j}\|\mathbf{x}^{i}-\mathbf{x}^{j}\|_{2}\geq\frac{1}{2DL}$, we have
$\frac{1}{T}\sum_{t\in[T]}\mathbb{P}_{t}(\mathbf{G}_{t})-\frac{1}{T}\sum_{t\in[T]}\mathbb{E}[\tilde{\ell}_{t}(\mathbf{G}^{*})]$ (20)
$\leq\mathcal{O}\Big{(}\frac{1}{T}\sqrt{M\sum_{t\in[T]}\sum_{j\in[M]}\mathbb{E}\big{\|}\mathbf{y}_{t}^{j}-\mathbf{y}_{t-1}^{j}\big{\|}_{2}^{2}}\Big{)}$
$\leq\mathcal{O}\Big{(}\frac{M}{T}\sqrt{\sum_{t\in[T]}\Big{[}\gamma_{X}^{2}\Big{(}\gamma_{H}^{2}-\mathbb{E}[\langle\mathbf{H}_{t},\mathbf{H}_{t-1}\rangle]\Big{)}+\sigma_{W}^{2}\Big{]}}\Big{)}.$
@@
Theorem 1 provides a performance guarantee of the learned channel decoder $\{\mathbf{G}_{t}\}_{t\in[T]}$: if $\exists\mathbf{G}^{*}\in\mathcal{G}$ such that the term $\sum_{t\in[T]}\mathbb{E}[\tilde{\ell}_{t}(\mathbf{G}^{*})]$ is sub-linear w.r.t $T$, i.e., $\sum_{t\in[T]}\tilde{\ell}_{t}(G^{*})=\mathcal{O}(T^{\alpha}),\alpha<1$, the average expected error probability satisfies $\frac{1}{T}\sum_{t\in[T]}\mathbb{P}_{t}(\mathbf{G}_{t})\rightarrow 0$ as $T\rightarrow\infty$. This is because $\mathbb{E}\|\mathbf{y}_{t}^{j}-\mathbf{y}_{t-1}^{j}\|_{2}^{2}\leq 4L^{2}$, and we have $\frac{1}{T}\sqrt{M\sum_{t\in[T]}\sum_{j\in[M]}\mathbb{E}\|\mathbf{y}_{t}^{j}-\mathbf{y}_{t-1}^{j}\|_{2}^{2}}\leq\frac{2ML\sqrt{T}}{T}=\mathcal{O}(\frac{\sqrt{T}}{T})$. Besides, we notice that a larger positive auto-correlation function $\mathbb{E}[\langle\mathbf{H}_{t},\mathbf{H}_{t-1}\rangle]$ of $\{\mathbf{H}\}_{t\in[T]}$ results in a tighter derived upper bound. This indicates that the proposed method utilizes the channel correlation (second-order channel statistics) to improve the performance of learned decoder.
@@
$\frac{1}{T}\sum_{t\in[T]}\mathbb{P}_{t}(\mathbf{G}_{t})-\frac{1}{T}\sum_{t\in[T]}\mathbb{E}[\tilde{\ell}_{t}(\mathbf{G}^{*})]$ (21)
$\leq\mathcal{O}\Big{(}\frac{M}{T}\sqrt{(\sigma^{\mathbf{H}}_{1:T})^{2}}+\frac{M}{T}\sqrt{(\Sigma^{\mathbf{H}}_{1:T})^{2}}+\frac{M\sqrt{T}}{T}\sqrt{\sigma_{W}^{2}}\Big{)},$
@@
Corollary 1 implies that the variance $\sigma^{2}_{W}$ of the channel noise $W$, the cumulative variance $(\sigma^{\mathbf{H}}_{1:T})^{2}$ and the cumulative variation $(\Sigma^{\mathbf{H}}_{1:T})^{2}$ of channel gain $\{\mathbf{H}_{t}\}_{t\in[T]}$ can deteriorate the performance of the learned channel decoder $\{\mathbf{G}_{t}\}_{t\in[T]}$. We observe that smaller $(\Sigma^{\mathbf{H}}_{1:T})^{2}$ leads to a tighter upper bound of average expected error probability $\frac{1}{T}\sum_{t\in[T]}\mathbb{P}_{t}(\mathbf{G}_{t})$, reflecting that the proposed method utilizes the distribution dependence of the time-correlated channel gain $\{\mathbf{H}_{t}\}_{t\in[T]}$ to enhance the online optimization process.
@@ V-A Preliminaries
One potential approach to tackle this challenge involves selecting codebooks for practical transmission from a pre-defined super-codebook $\mathbf{C}=\{C_{i}\}_{i=1}^{N}$ comprising $|\mathbf{C}|=N$ codebooks [8, 38]. This super-codebook can be statically constructed (such as a grid or a lattice [39]) in advance. To adapt to the changing statistical property of the channel noise, we dynamically choose one of these codebooks used for transmitting codewords in each round. Specifically, we select the codebook based on the symbol error rate of codebooks over the channel noise $Z_{t},\forall t\in[T]$. This approach draws inspiration from the Gibbs-algorithm-based codebook expurgation proposed in [8], and is amenable to practical deployment following the paradigm of Adaptive Modulation and Coding (AMC) widely deployed in communication systems [32]. Based on it, we consider that the super-codebook $\mathbf{C}$ is known to both the transmitter and receiver. Hence, the receiver is not required to transmit the entire learned codebook to the transmitter. Instead, the receiver only needs to send the index of the codebook selected in this round to the transmitter, thereby reducing the cost of executing online codebook learning.
@@ 1 Initialize: $\mathbf{w}^{\prime}_{1}=\arg\min_{\mathbf{w}\in\Omega}\psi_{1}(\mathbf{w})$ ;
$\frac{1}{T}\sum_{t\in[T]}\mathbb{P}_{t}(C_{t})$ (25)
$\quad+\frac{1}{T}\sqrt{\sum_{i\in[N]}\sum_{t\in[T]}\frac{1}{M}\sum_{j\in[M]}\sigma[\mathbbm{1}_{t}^{j}(i)]}\Big{)}+\frac{1}{T}\sum_{t\in[T]}\mathbb{E}[\ell_{t}(C^{*})]$
$\leq\tilde{\mathcal{O}}\Big{(}\frac{1}{T}\sqrt{N\sum_{t\in[T]}\sqrt{D_{kl}(\mathbb{P}_{t}\|\mathbb{P}_{t-1})}}$
$\quad+\frac{1}{T}\sqrt{\sum_{i\in[N]}\sum_{t\in[T]}\sum_{j\in[M]}\frac{\sigma[\mathbbm{1}_{t}^{j}(i)]}{M}}\Big{)}+\frac{1}{T}\sum_{t\in[T]}\mathbb{E}[\ell_{t}(C^{*})],$
where $\mathbbm{1}_{t}^{j}(i):=\mathbbm{1}\{\|\mathbf{x}_{t}^{j_{t}}-\mathbf{y}_{t}^{j}\|_{2}\leq\|\mathbf{x}_{t}^{j}-\mathbf{y}_{t}^{j}\|_{2}|\mathbf{a}_{t}=\mathbf{e}_{i}\}$ denotes the indicator of misclassifying the $j$-th codeword when the $i$-th codebook is selected in round $t$. $\mathbb{P}^{j}_{t}(i):=\mathbb{E}[\mathbbm{1}_{t}^{j}(i)]$ denotes the expectation of $\mathbbm{1}_{t}^{j}(i)$, and $\sigma[\mathbbm{1}_{t}^{j}(i)]:=\sqrt{\mathbb{E}[\mathbbm{1}_{t}^{j}(i)-\mathbb{P}^{j}_{t}(i)]^{2}}$ denotes the standard deviation of $\mathbbm{1}_{t}^{j}(i)$. $D_{kl}(\mathbb{P}_{t}\|\mathbb{P}_{t-1}):=\arg\max_{i\in[N],j\in[M]}D_{kl}(\mathbb{P}^{j}_{t}(i)\|\mathbb{P}^{j}_{t-1}(i))$ and $D_{kl}(P\|Q)=\mathbb{E}_{P}\log\frac{dP}{dQ}$ denotes the relative entropy between probability measures $P$ and $Q$.
@@
Theorem 2 presents theoretical performance guarantees of $\{C_{t}\}_{t\in[T]}$: if $\exists C^{*}\in\mathbf{C}$ such that $\sum_{t\in[T]}\ell_{t}(C^{*})$ is sub-linear w.r.t $T$, i.e., $\sum_{t\in[T]}\ell_{t}(C^{*})=\mathcal{O}(T^{\alpha}),\alpha<1$, then we have $\frac{1}{T}\sum_{t\in[T]}\mathbb{P}_{t}(C_{t})\rightarrow 0$, as $T\rightarrow\infty$. The reason is provided as follows. Since $1_{t}^{j}(i)$ is a Bernoulli random variable, we have $\sigma[1_{t}^{j}(i)]\leq 1$ and $|\mathbb{P}^{j}_{t}(i)-\mathbb{P}^{j}_{t-1}(i)|\leq 2$. Hence, we know $\frac{1}{T}\sqrt{\sum_{i\in[N]}\sum_{t\in[T]}\sum_{j\in[M]}\frac{\big{|}\mathbb{P}^{j}_{t}(i)-\mathbb{P}^{j}_{t-1}(i)\big{|}}{M}}\leq\sqrt{\frac{2N}{T}}$ and $\frac{1}{T}\sqrt{\sum_{i\in[N]}\sum_{t\in[T]}\sum_{j\in[M]}\frac{\sigma[1_{t}^{j}(i)]}{M}}\leq\sqrt{\frac{N}{T}}.$
In addition, $D_{kl}(\mathbb{P}_{t}\|\mathbb{P}_{t-1})$ essentially captures the discrepancy between the channel distributions of successive rounds. We notice that a smaller discrepancy $D_{kl}(\mathbb{P}_{t}\|\mathbb{P}_{t-1})$ results in a tighter regret bound. It indicates that we can learn the optimal codebook $C_{t}$ if the statistical properties of $Z_{t}$ remain relatively stable, implying that the proposed method leverages the distribution dependence of $\{Z_{t}\}_{t\in[T]}$ to improve the bandit learning process.